OpenStudy (anonymous):

x^2+3 over X ? I NEED HELP !

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

\[\frac{x^2+3}{x}\] is an expression what is the question?

OpenStudy (anonymous):

yess (:

OpenStudy (anonymous):

\[x+\frac{ 3 }{ x }\]

OpenStudy (misty1212):

lol is there a question?

OpenStudy (misty1212):

graph find the domain find the range plug in 2?

OpenStudy (anonymous):

fine the value exspression if X=3

OpenStudy (misty1212):

aaah i figured there was a question somewhere

OpenStudy (misty1212):

where you see an \(x\) replace it by a \(3\)

OpenStudy (misty1212):

\[\frac{\color{red}{x}^2+3}{\color{red}x}\] \[\frac{\color{red}{3}^2+3}{\color{red}3}\]

OpenStudy (anonymous):

\[\frac{ 3^{2}+3 }{ 3 }\]

OpenStudy (anonymous):

thank you <3

OpenStudy (anonymous):

now divide that and see what you get